The surface area of two cups are in the ratio 9:64 .if the smaller cup has a height of 25 cm and a volume of 2400 cm^3 calculate
a)the height of larger cup.
b)the volume of larger cup.

Answer by mananth(16946)

volume of cylinder = pir^2h
smaller cup
volume = 2400= pi*r^2*25
r^2=2400/(pi*25)
r=sqrt(2400/(pi*25)
r=5.3 cm
..
surface area of smaller cup = 2*pi*r*h
S=2*pi*5.3*25
S= 832.5 sq. cm
..
Ratio of surface area = 9/64
9/64= 832.5/x
x=64*832.5/9
= 5920 sq. cm is the surface area of larger cup